### v1

# Coupled Variational Reinforcement Learning for Language Model General Reasoning

Abstract
While reinforcement learning have achieved impressive progress in language model reasoning, they are constrained by the requirement for verifiable rewards. Recent verifier-free RL methods address this limitation by utilizing the intrinsic probabilities of LLMs generating reference answers as reward signals. However, these approaches typically sample reasoning traces conditioned only on the question. This design decouples reasoning-trace sampling from answer information, leading to inefficient exploration and incoherence between traces and final answers. In this paper, we propose C̱oupled V̱ariational Ṟeinforcement Ḻearning (CoVRL), which bridges variational inference and reinforcement learning by coupling prior and posterior distributions through a hybrid sampling strategy. By constructing and optimizing a composite distribution that integrates these two distributions, CoVRL enables efficient exploration while preserving strong thought-answer coherence. Extensive experiments on mathematical and general reasoning benchmarks show that CoVRL improves performance by 12.4% over the base model and achieves an additional 2.3% improvement over strong state-of-the-art verifier-free RL baselines, providing a principled framework for enhancing the general reasoning capabilities of language models.
Machine Learning, ICML
1 Introduction
Recent works (deepseekai2025deepseekr1incentivizingreasoningcapability; yue2025vapoefficientreliablereinforcement; yu2025dapoopensourcellmreinforcement) explore reinforcement learning with verifiable rewards (RLVR) to enhance LLMs’ reasoning capabilities, demonstrating impressive results in mathematical tasks. In this paradigm, the LLM generates a chain of thought (wei2023chainofthoughtpromptingelicitsreasoning) followed by an answer, and a rule-based program evaluates the final answer, assigning binary rewards based on correctness. The model is then optimized using policy gradient methods such as GRPO (shao2024deepseekmathpushinglimitsmathematical). However, these techniques require verifiable rewards, which restricts their applicability to other domains where formal verification of model answers is challenging or impossible.
Figure 1: Comparison between verifier-free RL with question-only training and CoVRL. Unlike prior methods that sample reasoning traces conditioned only on the question, CoVRL couples question-conditioned prior sampling with answer-conditioned posterior sampling via a hybrid variational framework, enabling efficient exploration while preserving strong trace–answer coherence.
A natural solution to this constraint is to introduce an LLM as a verifier (ma2025generalreasoneradvancingllmreasoning), which performs a role similar to that of the reward model in reinforcement learning from human feedback (ouyang2022traininglanguagemodelsfollow). However, this approach requires a strong verifier and faces the risk of reward hacking (skalse2025definingcharacterizingrewardhacking). As an alternative, verifier-free methods (yu2025rlprextrapolatingrlvrgeneral; tang2025verifiablerewardsscalingreinforcement; zhou2025reinforcinggeneralreasoningverifiers; chen2024languagemodelshiddenreasoners) have emerged that utilize the probabilities LLMs assign to correct answers as reward signals. These methods view answer generation as a process that relies on a set of implicit reasoning steps. In this formulation, each potential reasoning trace is treated as a latent variable contributing to the final answer. The probability of the correct answer is obtained by marginalizing over all plausible reasoning traces conditioned on the input question. In practice, the model is trained using reinforcement learning by sampling different intermediate thoughts for each question and using the probability of generating the correct answer from each thought as the reward.
While these methods eliminate external verifiers, they typically rely on question-only generation, i.e., the model observes the question and then attempts to generate reasoning steps leading to the final answer. As shown in Figure 1, this strategy faces two fundamental challenges: (1) Low sample efficiency, particularly for difficult questions where the model struggles to produce useful reasoning traces without guidance; and (2) Potential incoherence between reasoning traces and final answers due to the lack of answer guidance during trace generation, where even correct reasoning may receive low rewards because the final answer is expressed in a different format from the ground truth.
In this work, we propose Coupled Variational Reinforcement Learning (CoVRL), which frames reasoning training as a variational optimization problem that couples prior and posterior distributions. The core idea is to address the challenges of question-only training by establishing coupling between two complementary distributions for sampling reasoning traces during training. We leverage both prior and posterior distributions as variational components, corresponding to two complementary generation modes: (1) question-only generation, which samples traces from the prior distribution, reflecting real inference conditions; and (2) answer-guided generation, which samples traces from the posterior distribution to generate coherent reasoning that leads to the correct answer. This dual-mode strategy provides answer guidance during training while ensuring that the learned reasoning patterns transfer effectively to inference, thereby improving sample efficiency and mitigating incoherence between reasoning traces and answers.
To implement this, we construct a composite distribution that combines the prior and posterior probabilities, establishing a coupled training framework. Since directly sampling from this composite distribution is computationally complex, we adopt a hybrid sampling strategy where we randomly select between the two generation modes for each training example. We then maximize a variational lower bound, which includes a reconstruction term for answer prediction and a regularization term to ensure transferability to inference. Through importance weighting, we enable seamless training across both modes using the same underlying language model with different prompt templates.
In summary, our main contributions include:
  • We formulate reasoning optimization as a variational inference problem and introduce a composite distribution that theoretically unifies prior and posterior generation modes within a tractable framework.
  • We propose a hybrid sampling strategy that balances answer-guided learning and inference-time transferability, thus increasing sampling efficiency and addressing trace-answer coherence challenges.
  • We demonstrate consistent improvements across diverse benchmarks, achieving a 12.4% improvement over the base model and a 2.3% improvement over the strongest baseline.
2 Methodology
In this section, we introduce Coupled Variational Reinforcement Learning. Figure 2 illustrates the overall framework.
Figure 2: CoVRL employs hybrid sampling between prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ to generate reasoning traces. It optimizes the reconstruction term using GRPO and NLL loss, with KL regularization applied to ensure training-inference coherence.
2.1 Preliminary
Reinforcement Learning with Verifiable Rewards
Recent work (zhao2025geometricmeanpolicyoptimization; yu2025dapoopensourcellmreinforcement) has explored using reinforcement learning to optimize language models for reasoning tasks. Given a reward function $R(x,y)$ that evaluates the correctness of the generated reasoning and answer, the objective is:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},y\sim p_{\theta}(\cdot|x)}[R(x,y)]$ (1)
where $p_{\theta}$ denotes the model distribution parameterized by $\theta$, $x$ is the question, and $y$ represents the generated reasoning and answer. This objective can be optimized using policy gradient methods like GRPO (shao2024deepseekmathpushinglimitsmathematical). However, these approaches require verifiable reward signals, limiting their applicability to domains without accessible verifiers.
Variational Inference with Latent Variables
Variational inference (li2019stablevariationalautoencodertext; kingma2022autoencodingvariationalbayes) provides a principled framework for learning models with latent variables. Consider a generative process where observed data $y$ depends on latent variables $z$:
$p(y)=\int p(y|z)p(z)dz$ (2)
Since this marginal likelihood is typically intractable, we introduce a variational posterior $q_{\phi}(z|y)$. By applying Jensen’s inequality, we obtain:
$\log p(y)=\log\mathbb{E}_{q_{\phi}(z|y)}\left[\frac{p(y|z)p(z)}{q_{\phi}(z|y)}\right]$ (3)
$\geq\mathbb{E}_{q_{\phi}(z|y)}[\log p_{\theta}(y|z)]-D_{\text{KL}}(q_{\phi}(z|y)\|p(z))$
Here, $p_{\theta}(y|z)$ is the decoder distribution that reconstructs the observed data given the latent variables, while $q_{\phi}(z|y)$ is the encoder distribution that infers latent representations from the observed data. This lower bound is called the evidence lower bound (ELBO), which is widely used in variational autoencoders (kingma2022autoencodingvariationalbayes) and diffusion models (ho2020denoisingdiffusionprobabilisticmodels).
2.2 Coupled Variational Reinforcement Learning
In reasoning tasks, we treat reasoning traces as latent variables $z$ that mediate between questions $x$ and answers $y$, thus reformulating Equation 2 as:
$p(y|x)=\int p(y|z,x)p(z|x)dz$ (4)
Previous approaches generally optimize this objective by directly sampling from the prior distribution $p_{\phi}(z|x)$. However, these approaches face two main challenges: (1) Low sampling efficiency: reasoning traces sampled from the prior $p_{\phi}(z|x)$ may fail to effectively predict the target answer when the question is too difficult for the model, leading to inefficient exploration of the reasoning space. (2) Trace-answer incoherence: without access to the target answer, even correct reasoning traces can receive low rewards due to mismatched answer formulations or representations compared to the ground-truth answer $y$.
In response to these limitations, we introduce a variational distribution $q(z)$. Following the standard variational inference framework, we derive the following objective:
$\log p_{\theta}(y|x)=\log\int p_{\theta}(y|z,x)p_{\phi}(z|x)dz$ (5)
$=\log\int q(z)\frac{p_{\theta}(y|z,x)p_{\phi}(z|x)}{q(z)}dz$
$=\log\mathbb{E}_{q(z)}\left[\frac{p_{\theta}(y|z,x)p_{\phi}(z|x)}{q(z)}\right]$
$\geq\mathbb{E}_{q(z)}\left[\log\frac{p_{\theta}(y|z,x)p_{\phi}(z|x)}{q(z)}\right]\text{(Jensen's inequality)}$
$=\mathbb{E}_{q(z)}[\log p_{\theta}(y|z,x)]-D_{\text{KL}}(q(z)\|p_{\phi}(z|x))$
This ELBO consists of two components: a reconstruction term $\mathbb{E}_{q(z)}[\log p_{\theta}(y|z,x)]$ that encourages generating reasoning traces leading to the correct answer, and a KL regularization term $D_{\text{KL}}(q(z)\|p_{\phi}(z|x))$ that constrains deviation from the prior distribution.
Previous approaches can be viewed as setting $q(z)=p_{\phi}(z|x)$, which suffers from the aforementioned limitations. A natural alternative is to use the posterior distribution as $q(z)$, as in VAE (kingma2022autoencodingvariationalbayes), which samples reasoning traces conditioned on both the question and the target answer, i.e., $q(z)=q_{\psi}(z|x,y)$. This enables generating more relevant reasoning paths and improves sampling efficiency compared to prior-only sampling.
However, this introduces a fundamental training-inference mismatch. During training, we optimize the posterior distribution with access to target answers, but at inference time, we must rely on the prior distribution $p_{\phi}(z|x)$, which is conditioned only on the question. Even with KL divergence regularization $D_{\text{KL}}(q_{\psi}(z|x,y)\|p_{\phi}(z|x))$ that encourages the posterior to stay close to the prior, there is no guarantee that optimizing the posterior will improve the prior due to the asymmetric nature of reverse KL divergence (chan2022greedificationoperatorspolicyoptimization). Specifically, while $D_{\text{KL}}(q_{\psi}\|p_{\phi})$ constrains the posterior to avoid low-probability regions of the prior, it cannot ensure that the posterior covers all high-probability regions of the prior. This results in inadequate training for some regions of the prior, potentially leading to poor inference performance when sampling from these regions. In response to the difficulties of independently sampling from the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions, we introduce CoVRL, which establishes coupling between these distributions through complementary mechanisms.
2.3 Composite Distribution and Coupled Optimization
We first construct a composite distribution that unifies both question-only and answer-guided generation modes:
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$ (6)
This composite distribution establishes coupling between the prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions. By replacing $q(z)$ with this composite distribution $p^{\prime}(z|x,y)$, we reformulate the ELBO from Equation 5 as:
$\log p(y|x)=\log\int p(y|z,x)p(z|x)dz$ (7)
$=\log\mathbb{E}_{p^{\prime}(z|x,y)}\left[\frac{p(y|z,x)p(z|x)}{p^{\prime}(z|x,y)}\right]$
$\geq\underbrace{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{E}_{p^{\prime}(z|x,y)}[\log p_{\theta}(y|z,x)]}_{\text{Reconstruction term}}-\underbrace{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))}_{\text{Regularization term}}$
This ELBO consists of two components:
Reconstruction term
$\mathbb{E}_{p^{\prime}(z|x,y)}[\log p_{\theta}(y|z,x)]$ measures the model’s ability to predict the correct answer $y$ given the reasoning trace $z$ and question $x$, serving as the primary target for improving reasoning quality.
Regularization term
$D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ constrains the composite distribution to remain close to the prior distribution, ensuring that the learned reasoning patterns transfer effectively to inference.
This coupled optimization framework addresses the training-inference mismatch by simultaneously optimizing both the prior and posterior distributions toward the same objective through the composite distribution.
In practice, all three components can be implemented using a single LLM with different prompts: $p_{\phi}(z|x)$ conditions on the input question, $q_{\psi}(z|x,y)$ conditions on both question and answer, and $p_{\theta}(y|z,x)$ predicts the answer given the reasoning trace. However, directly sampling from the composite distribution $p^{\prime}(z|x,y)$ would require real-time mixing of prior and posterior distributions for generating each new token. This approach requires additional computation and complex modifications to prevalent LLM-RL frameworks (sheng2024hybridflow; hu2024openrlhf), which typically utilize efficient inference engines such as SGLang (sglang2024) and vLLM (kwon2023efficient) for rollout. Therefore, instead of sampling from the actual composite distribution, we adopt off-policy hybrid sampling. For each training sample, we randomly select from the prior with probability $\alpha$ or the posterior with probability $(1-\alpha)$:
$p_{\text{hybrid}}(z|x,y)=\begin{cases}p_{\phi}(z|x)&\text{with probability }\alpha\\ q_{\psi}(z|x,y)&\text{with probability }1-\alpha\end{cases}$ (8)
This approach ensures that both distributions are sampled, enabling joint optimization of exploratory and answer-guided reasoning.
For the Composite distribution
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$, the gradient becomes:
$\nabla_{\phi,\psi}\mathbb{E}_{p^{\prime}(z|x,y)}[$ $\log p_{\theta}(y|z,x)]=$ (9)
$\mathbb{E}_{p^{\prime}(z|x,y)}[$ $\log p_{\theta}(y|z,x)\nabla_{\phi,\psi}\log p^{\prime}(z|x,y)]$
Optimizing the composite distribution requires policy gradient methods due to the discrete sampling process. We employ Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical) to optimize the reconstruction term with the following objective:
[…]
where $r(z)=\frac{p^{\prime}_{\text{new}}(z|x,y)}{p^{\prime}_{\text{old}}(z|x,y)}$ represents the probability ratio between the new and old policies, and $\epsilon$ is the clipping parameter. The advantage $\hat{A}=\log p_{\theta_{\text{old}}}(y|z,x)-\bar{R}$ is computed using group relative estimation, where $\bar{R}$ is the average reward across the batch. Importantly, the reward is evaluated based on the model’s own answer prediction probability, eliminating the need for external verifiers.
However, the hybrid sampling strategy introduces a fundamental difference between the hybrid sampling distribution and the target composite distribution. Nevertheless, we can still optimize the composite distribution through importance sampling. Specifically, we replace $p^{\prime}_{\text{old}}(z|x,y)$ in Equation 10 with the actual sampling distribution $p_{\text{hybrid}}(z|x,y)$. Thus, the importance ratio $r(z)$ becomes:
$r(z)=\frac{p^{\prime}(z|x,y)}{p_{\text{hybrid}}(z|x,y)}$ (11)
This formulation enables mathematically principled optimization of the target composite distribution.
Prior distribution
For $p_{\phi}(z|x)$, the gradient is:
$\nabla_{\phi}$ $\mathbb{E}_{p_{\phi}(z|x)}[\log p_{\theta}(y|z,x)]=$ (12)
$\mathbb{E}_{p_{\phi}(z|x)}[\log p_{\theta}(y|z,x)\nabla_{\phi}\log p_{\phi}(z|x)]$
An important observation is that optimizing the prior distribution alone corresponds to standard maximum likelihood estimation, which can be efficiently computed using negative log-likelihood loss. However, applying NLL loss to all reasoning traces is problematic, as some traces may be of low quality. Since our model also serves as a reward model that evaluates reasoning quality through $\log p_{\theta}(y|z,x)$, we selectively reinforce high-quality traces. Therefore, we only compute the loss on samples that complete successfully and have positive advantage $\hat{A}>0$.
Prior Distribution Template <|im_start|>system A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think>reasoning process here</think><answer>answer here</answer>. <|im_end|> <|im_start|>user {question} <|im_end|> <|im_start|>assistant <think> {thought} </think> <answer> {answer} </answer> <|im_end|>
Posterior Distribution Template <|im_start|>system A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant provides the final answer first, then follows up with a comprehensive reasoning process. The answer and reasoning process are enclosed within <answer> </answer> and <think> </think> tags, respectively, i.e., <answer>answer here</answer><think>reasoning process here</think>. <|im_end|> <|im_start|>user {question} <|im_end|> <|im_start|>assistant <answer> {answer} </answer> <think> {thought} </think> <|im_end|>
Figure 3: Prompt templates after applying chat template for Prior and Posterior distributions. The key difference lies in the order of reasoning and answer components within the assistant’s response.
2.4 Optimizing KL Divergence
There are two approaches for optimizing the KL divergence term: (1) incorporating KL as part of the reward function as in PPO (schulman2017proximalpolicyoptimizationalgorithms), or (2) treating KL as a separate loss term as in GRPO (shao2024deepseekmathpushinglimitsmathematical). We adopt the latter approach, which we find is more stable.
Optimizing KL divergence as a loss function requires a non-negative and low-variance KL estimator. We extend existing KL estimators to handle our composite distribution setting, which couples prior and posterior through the KL regularization term. Specifically, we need to estimate $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ where $p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)$. We extend previous low-variance KL estimators (schulman2020kl) and derive different estimators depending on the sampling distribution using importance sampling with Bregman divergence-based control variates.
When sampling from the prior distribution:
$D_{\text{KL}}^{\text{prior}}=\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}\log\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}-\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}+1$ (13)
When sampling from the posterior distribution:
$D_{\text{KL}}^{\text{posterior}}=\frac{p^{\prime}(z|x,y)}{q_{\psi}(z|x,y)}\log\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}+\frac{p^{\prime}(z|x,y)}{q_{\psi}(z|x,y)}-1$ (14)
The key difference between these estimators lies in the sign of the Bregman correction term $\left(\frac{p^{\prime}(z|x,y)}{p_{\text{hybrid}}(z|x,y)}-1\right)$. This correction serves as a control variate that reduces estimation variance without introducing bias, leveraging the fact that:
$\mathbb{E}_{p_{\text{hybrid}}}[\frac{p^{\prime}(z|x,y)}{p_{\text{hybrid}}(z|x,y)}-1]=0$ (15)
The different signs ensure that the estimator remains non-negative regardless of whether the trace is sampled from the prior or posterior distribution. Moreover, we apply soft clipping to the importance sampling ratios in both estimators when they exceed a threshold to prevent numerical overflow from the exponential terms while still providing stable gradients. In cases where the model generation reaches the maximum response length, we skip computing the KL loss for these samples. Appendix A provides a more intuitive discussion of these KL estimators.
Table 1: Performance on benchmarks. Results are reported as Average@N where N indicates the number of independent evaluation runs to account for sampling variance. The Overall column presents the arithmetic mean across all benchmark tasks.
Method General Reasoning Mathematical Reasoning Overall
GPQA MMLU-Pro TheoremQA AIME’24 AQuA CARP-EN MATH-500 Minerva SAT-Math
Avg@4 Avg@1 Avg@2 Avg@32 Avg@4 Avg@2 Avg@4 Avg@4 Avg@32
Base Model 26.1 36.7 25.2 2.7 55.6 54.3 44.7 18.6 76.5 37.8
VeriFree 28.9 44.1 33.4 5.0 72.6 62.8 59.5 24.0 93.3 47.1
JLB 31.6 42.7 31.9 4.8 69.5 63.4 57.6 23.6 93.7 46.5
LaTRO 31.0 42.7 32.8 4.0 67.6 50.5 59.3 24.4 90.1 44.7
RAVR 27.6 34.2 24.9 3.4 59.9 53.4 44.0 13.2 81.2 38.0
RLPR 31.3 44.9 33.5 6.5 72.3 62.9 61.2 24.7 93.8 47.9
CoVRL (Ours) 30.4 46.5 36.3 7.5 77.3 65.1 66.3 25.5 97.1 50.2
2.5 Prompt Template and Tokenization
We follow the template format suggested in previous work (deepseekai2025deepseekr1incentivizingreasoningcapability), with simple extensions to suit our hybrid sampling strategy. The prompt templates for prior and posterior generation are shown in Figure 3. For tokenization, we separately encode special tokens including <think>, </think>, <answer> and </answer> to ensure consistent token ID sequences. However, during model generation, these special tokens may still be merged with adjacent tokens. Therefore, we use strings as stop signals for model generation, and after generation completes, we decode the outputs and re-encode them with separately encoded special tokens.
3 Experiment
3.1 Setup
Dataset
Following previous work, we use the dataset curated by (ma2025generalreasoneradvancingllmreasoning), which was sourced from WebInstruct (yue2024mammoth2scalinginstructionsweb). We retain only non-mathematical questions to assess training improvements on general reasoning abilities. Beyond this selection, we apply no additional filtering to evaluate the algorithm’s robustness across diverse question types, difficulty levels, and quality variations.
Implementation
We conduct experiments primarily using Qwen2.5-7B-Base (qwen2025qwen25technicalreport). We directly fine-tune the base model without an intermediate supervised fine-tuning stage. We implement our RL training pipeline using the verl framework (sheng2024hybridflow). During training, we use a batch size of 192 questions and generate 8 responses per question during rollout. For rollout sampling, we employ temperature=1.0, top_p=1.0, and max_tokens=2048 to ensure diverse response generation while maintaining reasonable response lengths. The clip threshold in the policy gradient loss is set to 0.3 given the off-policy nature of our algorithm. We use a cosine learning rate scheduler with 64 warmup steps and a peak learning rate of 1e-6.
Baselines
Our baseline selection focuses on other verifier-free methods that do not require external verification models during training: VeriFree (zhou2025reinforcinggeneralreasoningverifiers), RLPR (yu2025rlprextrapolatingrlvrgeneral), JLB (tang2025verifiablerewardsscalingreinforcement), RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning), and LaTRO (chen2024languagemodelshiddenreasoners). We summarize the gradient estimators of these methods in Appendix B. To eliminate the influence of different RL techniques and focus on the gradient estimator, we train all baselines on the Qwen2.5-7B-base model using the GRPO (shao2024deepseekmathpushinglimitsmathematical) with identical hyperparameters on the filtered dataset.
Evaluation
We evaluate our method on a comprehensive suite of reasoning benchmarks. For general reasoning, we use MMLU-Pro (wang2024mmluprorobustchallengingmultitask), GPQA (rein2023gpqagraduatelevelgoogleproofqa), and TheoremQA (chen2023theoremqatheoremdrivenquestionanswering). For mathematical reasoning, we use AIME’24 (aime2024_i), AQuA (ling2017programinductionrationalegeneration), CARP-EN (zhang2023evaluatingimprovingtoolaugmentedcomputationintensive), MATH-500 (lightman2023lets), Minerva (lewkowycz2022solvingquantitativereasoningproblems), and SAT-Math (mcaleste_sat_dataset_2023). For evaluation, we use temperature=0.6 and max_tokens=4096. We utilize Math-Verify to check answer correctness; more details are provided in Appendix C. To reduce evaluation variance, we repeat the evaluation $N$ times for each dataset, where $N$ depends on the test set size. We report the average performance across these $N$ runs as Average@$N$ to account for sampling variance.
(a) Reward Score
(b) Response Length
(c) NLL Loss
(d) KL Loss
Figure 4: Training dynamics of CoVRL across different metrics. We observe stable improvements in reasoning quality alongside effective optimization of both reconstruction and regularization objectives.
3.2 Main Results
Table 1 presents our main experimental results across all benchmarks. Our approach achieves consistent improvements over the base model and competitive performance against other verifier-free methods.
CoVRL achieves consistent performance improvements across benchmarks. CoVRL demonstrates substantial improvements over the base model across all tasks, achieving the highest overall performance at 50.2%, representing a 12.4% improvement. In comparison with other baselines, CoVRL achieves the strongest overall gains.
Reasoning capabilities learned by CoVRL are generalizable to mathematical reasoning. Despite training on non-mathematical questions, our approach shows substantial gains on mathematical benchmarks. This validates that general reasoning capabilities developed through diverse problem-solving can transfer effectively, highlighting the value of general reasoning skill development.
3.3 Training Dynamics
We present the evolution of key metrics in Figure 4 to better understand the algorithm. This reveals several observations:
The posterior distribution is effective in providing guidance.
Figure 4(a) shows steady reward score improvements throughout training, with the posterior distribution consistently outperforming the prior distribution. This persistent score gap validates our answer-guided sampling strategy and confirms that posterior sampling enables more efficient exploration of high-quality reasoning paths.
CoVRL enhances reasoning capabilities with prolonged chain-of-thought traces.
Figure 4(b) demonstrates a stable increase in response length, indicating that the model progressively generates more detailed reasoning processes. This trend suggests that CoVRL successfully encourages elaborate chain-of-thought reasoning with more thorough step-by-step explanations.
Regularization provides stable optimization dynamics.
Figures 4(c) and 4(d) show stable downward trends in NLL and KL losses. The decreasing losses indicate improved answer prediction and successful regularization, confirming that our variational objective effectively balances the reconstruction and regularization terms.
Figure 5: Impact of prior sampling probability $\alpha$ in hybrid sampling strategy.
3.4 Hybrid Sampling Strategy
We examine the impact of our hybrid sampling strategy by varying the mixing ratio $\alpha$ between prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions. Here $\alpha$ represents the probability of sampling from the prior distribution. By default, we set $\alpha=0.5$. Figure 5 demonstrates that low prior sampling probability ($\alpha=0.1$) outperforms high prior sampling probability ($\alpha=0.9$). This highlights the important role of the posterior distribution in our algorithm. When $\alpha=0.9$, the model primarily samples from the prior distribution, and we observe reduced reasoning chain length. We attribute this to difficulty in improving rewards, leading the model to prioritize minimizing KL loss and generating shorter sequences. When $\alpha=0.1$, reasoning chain length increases, and posterior-dominated sampling achieves better performance than prior-dominated sampling. However, due to training-inference mismatch, performance remains inferior to balanced sampling. Training dynamics under these settings are provided in Appendix D.
Table 2: Performance comparison across different base models of various model architectures and sizes.
General Mathematical
Model GPQA MMLU-Pro TheoremQA AIME’24 AQuA CARP-EN MATH-500 Minerva SAT-Math Overall
Avg@4 Avg@1 Avg@2 Avg@16 Avg@4 Avg@2 Avg@4 Avg@4 Avg@32
Qwen2.5-7B 26.1 36.7 25.2 2.7 55.6 54.3 44.7 18.6 76.5 37.8
   + CoVRL 30.4+4.3 46.5+9.8 36.3+11.1 7.5+4.8 77.3+21.7 65.1+10.8 66.3+21.6 25.5+6.9 97.1+20.6 50.2+12.4
Qwen2.5-14B 28.0 38.5 25.4 3.0 63.3 57.3 44.8 22.1 82.5 40.5
   + CoVRL 36.6+8.6 57.0+18.5 42.4+17.0 7.9+4.9 81.8+18.5 64.0+6.7 71.5+26.7 33.7+11.6 95.8+13.3 54.5+14.0
Qwen3-8B 36.9 51.0 32.8 4.0 68.6 59.4 53.6 30.0 90.5 47.4
   + CoVRL 37.6+0.7 59.6+8.6 43.7+10.9 9.2+5.2 80.5+11.9 66.1+6.7 74.5+20.9 35.2+5.2 97.6+7.1 56.0+8.6
Qwen3-14B 37.5 57.6 37.6 7.8 75.3 64.6 67.2 33.5 93.1 52.7
   + CoVRL 42.7+5.2 63.1+5.5 46.3+8.7 9.5+1.7 83.9+8.6 65.9+1.3 76.1+8.9 38.1+4.6 97.4+4.3 58.1+5.4
Table 3: Performance comparison across different training data compositions.
General Mathematical
Training Data GPQA MMLU-Pro TheoremQA AIME’24 AQuA CARP-EN MATH-500 Minerva SAT-Math Overall
Avg@4 Avg@1 Avg@2 Avg@16 Avg@4 Avg@2 Avg@4 Avg@4 Avg@32
Base Model 26.1 36.7 25.2 2.7 55.6 54.3 44.7 18.6 76.5 37.8
Non-Math Only 30.4+4.3 46.5+9.8 36.3+11.1 7.5+4.8 77.3+21.7 65.1+10.8 66.3+21.6 25.5+6.9 97.1+20.6 50.2+12.4
Math Only 28.9+2.8 42.7+6.0 31.8+6.6 4.2+1.5 72.1+16.5 53.1-1.2 60.1+15.4 20.6+2.0 94.3+17.8 45.3+7.5
3.5 Training Model and Data
CoVRL is robust across different base models.
We evaluate CoVRL on Qwen2.5-base and Qwen3-base models ranging from 7B to 14B parameters. Note that for Qwen3-base experiments, we replace <think> </think> with <thinking> </thinking> tags, as Qwen3-base fails to properly utilize <think> </think> tags in prompts while responding well to other similar tags. As shown in Table 2, CoVRL delivers consistent performance improvements across all tested models, demonstrating its robustness across model architectures.
CoVRL learns generalizable reasoning abilities from training data.
We also evaluate CoVRL with different training data compositions. The results in Table 3 reveal that training on both math-only and non-math-only data significantly improves base model performance. Notably, models trained solely on mathematical data demonstrate enhanced performance on non-mathematical reasoning tasks, and similarly, models trained on non-mathematical data improve on mathematical tasks. This indicates that our method enables models to acquire generalizable reasoning capabilities that transfer across different domains.
4 Related Works
Verifier-free Reinforcement Learning
While reinforcement learning from verifiable rewards (cui2025processreinforcementimplicitrewards; yu2025dapoopensourcellmreinforcement; yang2025qwen3technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability) has emerged as a common practice, these approaches are restricted to domains where robust verifiers are available. Recent work explores eliminating external verification dependencies. LaTRO (chen2024languagemodelshiddenreasoners) proposes a framework that formulates reasoning as sampling from a latent distribution. JLB (tang2025verifiablerewardsscalingreinforcement) attempts to scale this approach to data that cannot be easily verified. VeriFree (zhou2025reinforcinggeneralreasoningverifiers) shares a similar framework and proposes using probability directly as both a reward and a coefficient for negative log-likelihood loss. RLPR (yu2025rlprextrapolatingrlvrgeneral) introduces a different reward mechanism using the mean of token rewards. These approaches sample from prior distributions, potentially suffering from limited sample efficiency and reasoning-answer misalignment issues. In contrast, RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning) also utilizes posterior distributions. Compared to RAVR, our approach establishes explicit coupling between prior and posterior through a composite distribution and hybrid sampling strategy.
Self-improving language models
Recent work explores various strategies for enhancing language models by leveraging their own outputs. One line focuses on self-generated training signals, including methods that prompt models to rank their responses for preference learning (yuan2025selfrewardinglanguagemodels), apply iterative DPO (chen2025bootstrappinglanguagemodelsdpo) with implicit rewards (rafailov2024directpreferenceoptimizationlanguage), and employ majority voting for test-time improvements (zuo2025ttrltesttimereinforcementlearning). Alternative approaches construct rewards based on reference answers, such as the verifier-free methods discussed above. Among them, zhou2025variationalreasoninglanguagemodels also introduce posterior distributions and update two separate prior and posterior models using the IWAE paradigm (burda2016importanceweightedautoencoders) with an EM algorithm, utilizing external verifiers for variance reduction. In comparison, we adopt online RL training to optimize a single composite distribution with verifier-free rewards.
5 Conclusion
In this paper, we introduce Coupled Variational Reinforcement Learning, which enables joint optimization of question-only and answer-guided generation while maintaining theoretical grounding through a coupled optimization framework based on composite distributions. Experimental results demonstrate consistent improvements across tasks, validating the effectiveness of our approach.
Appendix A KL Divergence Estimator
In this section, we provide an intuitive understanding of our KL divergence estimators and derive the specific forms used in our hybrid sampling strategy.
Let $r=\frac{q_{\psi}(z|x,y)}{p_{\phi}(z|x)}$ denote the likelihood ratio between posterior and prior distributions. Our composite distribution becomes:
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)=p_{\phi}(z|x)\left(\frac{1}{2}+\frac{1}{2}r\right)$
We need to estimate $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ using importance sampling from either the prior or posterior distribution.
When sampling from the prior distribution:
The importance sampling ratio is $\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}=\frac{1}{2}+\frac{1}{2}r$, and our KL estimator becomes:
$D_{\text{KL}}^{\text{prior}}=\left(0.5+0.5r\right)\ln\left(0.5+0.5r\right)-\left(0.5r-0.5\right)$
When sampling from the posterior distribution:
The importance sampling ratio is $\frac{p^{\prime}(z|x,y)}{q_{\psi}(z|x,y)}=\frac{0.5+0.5r}{r}=\frac{0.5}{r}+0.5$, and our KL estimator becomes:
$D_{\text{KL}}^{\text{posterior}}=\left(\frac{0.5}{r}+0.5\right)\ln\left(0.5+0.5r\right)+\left(\frac{0.5}{r}-0.5\right)$
Figure 6 shows the behavior of both estimators across different likelihood ratios. When $r>1$, the posterior assigns higher probability to the sample than the prior, indicating reasoning traces well-aligned with target answers. When $0<r<1$, the prior assigns higher probability, representing more exploratory reasoning paths. Our hybrid sampling strategy leverages this complementary behavior to ensure stable optimization across diverse reasoning patterns.
Figure 6: Comparison of KL divergence estimators under different sampling strategies.
Appendix B Comparison between Existing Approaches
We compare the gradient estimators of various verifier-free reasoning methods:
$\nabla_{\theta}J_{\text{JLB}}$ $=\mathbb{E}_{z\sim\pi_{\theta}(z|x)}\big[\log\pi_{\theta}(y^{*}|x,z)\nabla_{\theta}\log\pi_{\theta}(z|x)+1\cdot\nabla_{\theta}\log\pi_{\theta}(y^{*}|x,z)\big]$ (16)
$\nabla_{\theta}J_{\text{LaTRO}}$ $=\mathbb{E}_{z\sim\pi_{\theta}(z|x)}\big[(\log\pi_{\theta}(y^{*}|x,z)-\beta\log\frac{\pi_{\theta}(z|x)}{\pi_{\text{ref}}(z|x)})\nabla_{\theta}\log\pi_{\theta}(z|x)+1\cdot\nabla_{\theta}\log\pi_{\theta}(y^{*}|x,z)\big]$
$\nabla_{\theta}J_{\text{VeriFree}}$ $=\mathbb{E}_{z\sim\pi_{\theta}(z|x)}\big[\pi_{\theta}(y^{*}|z,x)\nabla_{\theta}\log\pi_{\theta}(z|x)+\pi_{\theta}(y^{*}|z,x)\nabla_{\theta}\log\pi_{\theta}(y^{*}|x,z)\big]$
$\nabla_{\theta}J_{\text{RLPR}}$ $=\mathbb{E}_{z\sim\pi_{\theta}(z|x)}\big[\frac{\sum(\{p_{i}|t_{i}\in y^{*}\})}{|y^{*}|}\nabla_{\theta}\log\pi_{\theta}(y^{*},z|x)\big]$
$\nabla_{\theta}J_{\text{RAVR}}$ $=\mathbb{E}_{z\sim\pi_{\theta}(z|x,y^{*})}\big[R(z)\nabla_{\theta}\log\pi_{\theta}(z|x)\big]+\nabla_{\theta}R(z)\times D_{\text{KL}}(p(z|x,y^{*})\|p(z|x))\text{, where}$
$R(z)$ $=\max(0,\log\pi_{\theta}(y^{*}|z,x)-\mathbb{E}_{z^{\prime}\sim\pi_{\theta}(z|x)}[\log\pi_{\theta}(y^{*}|x,z^{\prime})])$
All methods aim to improve reasoning without external verifiers but differ in their approach. JLB (tang2025verifiablerewardsscalingreinforcement) uses log-probability as reward with fixed answer term weighting. LaTRO (chen2024languagemodelshiddenreasoners) incorporates KL regularization between policy and reference models. VeriFree (zhou2025reinforcinggeneralreasoningverifiers) uses the probability instead of log-probability as reward and also uses it as the weight term for NLL loss. RLPR (yu2025rlprextrapolatingrlvrgeneral) optimizes joint answer-reasoning probability using token-level probabilities as rewards. RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning) samples from posterior distributions and optimizes the corresponding variational lower bound. The core differences among these methods lie in sampling distribution, reward design, and the weighting coefficients for the NLL loss terms.
In contrast, our CoVRL introduces a fundamentally different approach through hybrid sampling and composite distribution optimization. The gradient combines three key components: (1) importance-weighted policy gradient on the composite distribution $p^{\prime}(z|x,y)$ for efficient exploration-exploitation balance, (2) selective NLL loss applied only to high-quality samples ($\mathbb{I}[\hat{A}>0]$), and (3) explicit KL regularization to ensure transferability from training to inference.
$\nabla J_{\text{CoVRL}}$ $=\mathbb{E}_{z\sim p_{\text{hybrid}}(z|x,y)}\big[\frac{p^{\prime}(z|x,y^{*})}{p_{\text{hybrid}}(z|x,y^{*})}\times\log{p_{\theta}(y^{*}|x,z)}\nabla\log p^{\prime}(z|x,y)$ (17)
$\quad+\mathbb{I}[\hat{A}>0]\nabla\log p_{\theta}(y^{*}|x,z)\big]-\nabla D_{\text{KL}}(p^{\prime}(z|x,y^{*})\|p_{\phi}(z|x))$
CoVRL leverages both question-only and answer-guided sampling through $p_{\text{hybrid}}(z|x,y)$, enabling more effective training while maintaining inference compatibility.
Appendix C Evaluation Implementation
For non-multiple-choice questions, we use the math-verify library to directly parse the standard answers. For multiple-choice questions, we standardize all datasets into a unified multiple-choice format. When parsing model responses, we employ different strategies based on the question type:
Non-multiple-choice questions:
We first use the default parsing method of the math-verify library to extract results. If no parsing result is obtained, we wrap the answer with \$\$ to parse it as a LaTeX expression. This approach handles cases where the model outputs <answer>latex expression</answer> without proper LaTeX delimiters. If there is still no valid parsing result, we perform exact string matching as a fallback.
Multiple-choice questions:
When parsing standard answers, we include both the option labels (e.g., A, B, C, D) and the option content as potential matches, since the model sometimes outputs the option content instead of the option label. When parsing model outputs, we first attempt to match the option format. If option matching fails, we try to parse the response as LaTeX format. If LaTeX parsing also fails, we perform exact string matching detection.
(a) Prior Reward Score
(b) Posterior Reward Score
(c) Response Length
Figure 7: Training dynamics comparison between different prior sampling probabilities ($\alpha=0.1$ vs $\alpha=0.9$).
Appendix D Training Dynamics of Different Sampling Ratios
Figure 7 illustrates the distinct training dynamics across different sampling ratios. As shown in the figures, $\alpha=0.9$ achieves higher prior reward scores, while $\alpha=0.1$ obtains higher posterior reward scores. This aligns with expectations, as models trained with more prior sampling naturally perform better under prior evaluation conditions, while those trained with more posterior sampling excel under posterior evaluation conditions. Additionally, response lengths increase with $\alpha=0.1$ but decrease with $\alpha=0.9$ during training. This shows that posterior-dominated sampling encourages longer reasoning chains guided by answer information, while prior-dominated sampling leads to shorter responses due to stronger KL regularization pressure. When prior sampling dominates, the model struggles to improve reconstruction rewards and instead favors minimizing KL divergence by generating shorter responses. In contrast, posterior guidance enables consistent reward improvements that incentivize more detailed reasoning. These dynamics provide additional evidence for the effectiveness of balanced sampling strategies.
Appendix E Influence of Loss Components
We conduct comprehensive ablation studies to understand the impact of key components in our CoVRL framework, including regularization coefficients and reward function formulations. The results are presented in Table 4.
Table 4: Ablation study on regularization coefficients and reward functions.
General Mathematical
Configuration GPQA MMLU-Pro TheoremQA AIME’24 AQuA CARP-EN MATH-500 Minerva SAT-Math Overall
Avg@4 Avg@1 Avg@2 Avg@16 Avg@4 Avg@2 Avg@4 Avg@4 Avg@32
Regularization Coefficients
Default 30.4 46.5 36.3 7.5 77.3 65.1 66.3 25.5 97.1 50.2
$\lambda_{KL}=0.1$ 31.3 32.5 20.7 0.3 35.1 33.8 18.1 11.9 62.6 27.4
$\lambda_{NLL}=0.1$ 31.2 42.5 33.1 4.2 69.0 49.6 59.5 23.5 89.6 44.7
Reward Functions
Log-Probability 30.4 46.5 36.3 7.5 77.3 65.1 66.3 25.5 97.1 50.2
Probability 31.4 46.9 35.8 6.3 78.2 65.3 67.6 27.0 94.9 50.4
Probability Sum 28.1 45.9 37.2 7.3 74.8 65.0 68.7 24.9 95.1 49.7
Log-Probability Sum 31.5 47.7 36.7 7.1 76.8 65.1 67.1 27.4 96.0 50.6
KL regularization
The results show that reducing the KL divergence coefficient to 0.1 significantly degrades performance across all benchmarks, with overall accuracy dropping to 27.4%. This performance degradation stems from training instability when the KL regularization is insufficient. We observe substantial increases in KL divergence during training, indicating significant deviation between the prior and posterior distributions. This leads to both training-inference mismatch issues and training instability, as we are essentially performing off-policy optimization with an increasing distribution shift between training and inference.
NLL loss
In contrast, the model appears less sensitive to changes in the NLL loss coefficient. When reducing the NLL coefficient to 0.1, performance decreases moderately to 44.7%. We attribute this resilience to the fact that the RL term and NLL loss optimize essentially the same objective, both of which aim to improve answer prediction quality. The NLL loss primarily trains the model’s ability to summarize reasoning and produce final answers.
Reward function formulation
We examine different reward formulations for our variational framework, specifically focusing on (1) Length normalization: comparing averaging over sequence length versus unnormalized probability sums; and (2) Logarithmic transformation: examining whether to use log-probabilities or raw probabilities as reward signals. The results demonstrate that all reward formulations achieve remarkably similar overall performance, with variations of less than 1 percentage point (49.7% to 50.6%). This consistency indicates that our CoVRL framework is robust to various reward formulations.